In Young’s double slit experiment, the fringe width is ß. If the entire arrangement is placed in a liquid of refractive index n, the
fringe width becomes :
(A) nß (B) +1
ß
n (C) -1
ß
n (D) n
ß

Grade:

2 Answers

Arun
3 years ago
If the screen distance id D, distance between the slits is d and the wave length is λ
then the fringe width is given by
b = λD/d
When the arrangement is immerse in a liquid of refractive index u then the new wavelength is 1/u times the original wavelength
If the new wavelength is λ' then
λ' = λ/n
Therefore, the new fringe width will be
b' = λ'D/d
or, b' = b/u
Therefore the fringe width will decrease by a factor of u.
 
Vikas TU
3 years ago
Dear student 
Question is not clear 
Please attach an image, 
We will happy to  help you 
Good Luck
Cheers